OpenStudy (anonymous):

what are the real or imaginary solutions of the polynomial equation x^3-8

sam (.sam.):

Let \(p(x)=x^3-8\) find zeros when \(p(x)=0\) \[x^3-8=0 \\ \\ x^3=8 \\ \\ x=2\] 1 real solution, can you see it now?

OpenStudy (raden):

hint : x^3 - 8 = (x-2)(x^2+2x+4)

OpenStudy (raden):

then make all factors be equal zero, solve for x

OpenStudy (anonymous):

@.Sam. : It is x^3 NOT x^2

sam (.sam.):

The other 2 are imaginary

OpenStudy (anonymous):

@.Sam. : What do you mean? The solution shall be x=-2 and x=+2 (with +2 being repeated roots). So yeah, it has 2 solutions +2 and -2, but your's isn't correct.

sam (.sam.):

The real solution of p(x) is x=2, (x-2) \(is \) a factor of p(x), Then you'll use (x-2) do synthetic division and get \(x^2+2x+4\) \[=(x-2)(x^2+2x+4)\] Then just find the imaginary solution on \(x^2+2x+4\)

OpenStudy (reemii):

you must solve (in the complex plane) the equation \(z^3 = 8\), that is, in polar coordinates \(r^3e^{i3\theta} = 2\). you'll get 3 solutions, and one of those is \(z=2\).

OpenStudy (anonymous):

x^2+2x+4= (x+2)^2... the roots of this is x=-2. -2 is a real number. There is no complex solution to it. @reemii - Yeah, one can do it as well, given all real numbers can be represented on the complex plane.

OpenStudy (anonymous):

umm... so it's not 2?

OpenStudy (raden):

well to get roots of x^2 + 2x + 4 = 0, use the quadratic formula : x = (-b +- sqrt(b^2 - 4ac))/(2a) here, given a = 1, b=2 and c = 4

OpenStudy (anonymous):

alright, thank you very much.

OpenStudy (raden):

and actually the solutions above be complex, because the value of discriminant (D) < 0 with D = b^2 - 4ac